Circular No. 6509

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-3 W. Cui, Massachusetts Institute of Technology (MIT), reports for the

RXTE ASM Team at MIT and the Goddard Space Flight Center: “Cyg X-3 has just started another x-ray outburst, as seen by the RXTE All Sky Monitor. This is only ∼ 3 weeks after the end of the previous one, which peaked at ∼ 220 mCrab (1.3–12 keV) and lasted for ∼ 50 days

(see the ASM light curves at http://space.mit.edu/XTE/ASM lc.html).

The daily averaged ASM measurements show that it has already reached a level of ∼ 300 mCrab (299 ± 7 mCrab on Nov. 21), which is about a factor of 3 higher than its ‘normal’ persistent flux level. During this active period, observations of this source at other wavelengths, especially radio and infrared, are strongly encouraged.”

SUPERNOVA 1996bu IN NGC 3631 M. Armstrong, Rolvenden, U.K.; and G. M. Hurst, The Astronomer,

report that Armstrong recorded a prediscovery image of SN 1996bu (cf. IAUC 6505) on an unfiltered CCD frame obtained with a 0.26-m reflector on Nov. 14.212 UT for the U.K. Nova/Supernova Patrol; the supernova’s magnitude was estimated as 18.0.

COMET C/1996 E1 (NEAT)

Extension to the ephemeris on IAUC 6481:

1996/97 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 23 20^{h}36.^{m}27 + 3^{o}15.^{′}2 2.227 2.150 72.^{o}7 26.^{o}0 15.4
Dec. 3 20 47.99 − 0 11.1 2.493 2.252 64.5 23.3 16.0
13 20 59.37 − 2 44.9 2.754 2.355 56.3 20.4 16.6
23 21 10.46 − 4 40.3 3.003 2.458 48.1 17.3 17.2
Jan. 2 21 21.27 − 6 07.5 3.237 2.562 39.9 14.3 17.7
12 21 31.81 − 7 13.7 3.452 2.666 31.8 11.2 18.2

Total visual magnitude estimates: Oct. 9.16 UT, 13.1 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 13.80, 13.5 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector); Nov. 9.79, 14.6 (Sárneczky).

1996 November 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green